# InternLM-XComposer2.5-Reward: A Simple Yet Effective Multi-Modal Reward Model

### v1

@@ 5 Experiments
Table 4: Evaluation results on Reward Bench [41]. We report the performance of selective representative language-only RMs and previous multi-modal generative RMs.
@@
InternLM2-7B-Reward [8] InternLM2-7B 99.2 69.5 87.2 94.5 87.6
InternLM2-20B-Reward [8] InternLM2-20B 98.9 76.5 89.5 95.8 90.2
Skyword-Reward-Llama3.1-8B [48] Llama3.1-8B 95.8 87.3 90.8 96.2 92.5
INF-ORM-Llama3.1-70B [97] Llama3.1-70B 96.6 91.0 93.6 99.1 95.1
@@
QWen2-VL-7B [90] QWen2-7B 96.6 57.0 73.9 94.3 83.8
LLaVA-Critic-8B [94] LLaMA3-7B 96.9 52.8 81.7 83.5 80.0
IXC-2.5-Reward (Ours) InternL2-7B 90.8 83.8 87.8 90.0 88.6
Table 5: Evaluation results on RM-Bench [51]. We classify reward models into three types: sequence classifiers (Seq.), generative models, and implicit DPO models. Performance is reported across four domains (Chat, Math, Code, Safety) and three difficulty levels (Easy, Normal, Hard), along with average scores (Avg).
@@
Tulu-2-dpo-13b [32] Implicit 66.4 51.4 51.8 85.4 86.9 66.7 37.7 63.8
InternLM2-7B-Reward [8] Seq. 61.7 71.4 49.7 85.5 85.4 70.7 45.1 67.1
InternLM2-20B-Reward [8] Seq. 63.1 66.8 56.7 86.5 82.6 71.6 50.7 68.3
Nemotron-4-340B-Reward [92] Generative 71.2 59.8 59.4 87.5 81.0 71.4 56.1 69.5
URM-LLaMa-3.1-8B [55] Seq. 71.2 61.8 54.1 93.1 84.0 73.2 53.0 70.0
Skyword-Reward-Llama3.1-8B [48] Seq. 69.5 60.6 54.5 95.7 89.0 74.7 46.6 70.1
@@
Instruction WildVision(0617) [62] Open 89.2 [31] 67.3 [94] 37.5 74.6
Following MIA${}_{(\text{val})}$ [68] Open 88.6 [31] 80.7 [90] 80.4 84.0
& Chat MM-MT${}_{(\text{val})}$ [1] Open 7.72 [31] 5.45 [90] 3.85 5.70
MM-Vet v2${}_{(\text{0613})}$ [103] Open 71.8 [4] 58.1 [14] 45.8 54.8
Knowledge MMBench${}_{(\text{v1.1})}$ [52] MCQ 85.7 [74] 82.7 [61] 79.4 79.0
MMMU${}_{(\text{val})}$ [106] MCQ 70.7 [31] 56.2 [14] 42.9 44.1
MMStar [12] MCQ 72.7 [74] 63.2 [14] 59.9 59.6
Reasoning MathVista${}_{(\text{mini})}$ [59] VQA 78.4 [74] 66.5 [60] 63.7 63.4
MathVerse${}_{(\text{vision-only})}$ [113] VQA 54.8 [26] 26.6 [50] 16.2 19.0
MathVision${}_{(\text{full})}$ [89] VQA 43.6 [26] 22.0 [50] 17.8 18.8
Text-Rich TextVQA${}_{(\text{val})}$ [79] VQA 82.0 [64] 78.5 [42] 78.2 81.3
ChartQA${}_{(\text{test})}$ [63] VQA 81.2 [64] 82.4 [99] 82.2 80.5
OCRBench [49] VQA 89.4 [74] 82.2 [14] 69.0 70.0
@@
5.1 Evaluation Results of IXC-2.5-Reward
Benchmarks
To evaluate IXC-2.5-Reward, we use diverse reward model benchmarks: (1) VL-RewardBench [43], encompassing 1250 multi-modal problems addressing general understanding, hallucination, and reasoning challenges; (2) Reward-Bench [41], with 2985 language-only problems including chat, chat hard, safety and reasoning; and (3) RM-Bench [51], comprising 1237 language-only problems across chat, math, code, and safety. RM-Bench defines three tracks (easy, normal, hard) that evaluate the sensitivity of reward models to subtle content variations and style biases. While Reward-Bench and RM-Bench are designed for reward models of language-only LLMs, we evaluate IXC-2.5-Reward on these benchmarks to demonstrate that our multi-modal reward model maintains strong language capabilities despite also processing image and video inputs.
5.1.1 Results on VL-RewardBench
Main Results
Tab. 3 presents the evaluation results of various multi-modal RMs on the VL-RewardBench [43]. Unlike previous multi-modal generative reward models, our IXC-2.5-Reward is a discriminative model that predicts a scalar reward. Our proposed IXC-2.5-Reward model, despite being an open-source 7B parameter model, outperforms all other open-source models. Notably, IXC-2.5-Reward achieves the highest overall accuracy (65.8%) among open-source models and the highest Macro Accuracy (70.0%) among all models, indicating its superior performance in handling diverse tasks within the VL-RewardBench.
Strong Performance on General Problems
The results in Table 3 reveal that IXC-2.5-Reward achieves a significantly higher accuracy (84.7%) on general problems compared to other generative RMs. We found the reason is attributed to these problems presenting a considerable challenge, often leading to tied judgments in previous LVLMs, whereas IXC-2.5-Reward demonstrates a greater ability to make correct classifications with different scalar scores.
5.1.2 Results on Reward Bench and RM-Bench
Main Results
We argue that multi-modal RMs should preserve strong language processing abilities despite the incorporation of image and video data during training. Consequently, we evaluate the performance of multi-modal reward models, including IXC-2.5-Reward, on Reward Bench (Tab. 4) and RM-Bench (Tab. 5). The results demonstrate that IXC-2.5-Reward achieves considerable performance and surpasses other multi-modal models on this benchmark.
Sensitivity to Content and Style
Consistent with findings in [51], IXC-2.5-Reward demonstrates sensitivity to subtle content variations and style biases, an issue often overlooked in multi-modal research. We believe further research is needed to enhance the robustness of multi-modal reward models.
5.2 Evaluation Results of IXC-2.5-Chat
Benchmarks
We select four representative benchmarks for evaluating the instruction following and in-the-wild chatting abilities of LVLMs. (1) The WildVision bench [62] uses prompts collected from user submissions, reflecting real-world multimodal interactions. (2) MIA-bench [68] that is specially designed to evaluate instruction following. (3) MM-MT [1] which is an instruction-following benchmark for multi-modal models, exhibits a strong correlation with LMSys-Vision ELO ratings [15]. (4) MM-Vet [103] that evaluate LVLMs on complex tasks such as language generation. These datasets contain open-ended questions and the referenced answers, and evaluation is performed using an LLM-as-a-Judge [118] approach, which involves using a judge model like GPT-4o [31] to predict scores.
We also report the performance on other categories, such as MMBench [52], MMMU [106] and MMStar [12] for general knowledge, MathVerse [113] and MathVision [89] for math reasoning, TextVQA [79], ChartQA [63] and OCRbench [49] for text-rich document understanding. These benchmarks utilize multiple-choice questions (MCQ) or visual question answering (VQA), where responses are limited to short keywords and evaluated based on string matching.
Results on Instruction Following & Chat
Tab. 6 shows that IXC-2.5-Chat outperforms previous SOTA models across multiple benchmarks (WildVision, MIA, and MM-MT), demonstrating significant improvements in multi-modal understanding with instruction following ability and provide more comprehensive information for in-the-wild chat scenario.
Results on Other Categories
On other categories (Knowledge, Reasoning, and Text-Rich), IXC-2.5-Chat performs comparably to the supervised fine-tuned (SFT) model IXC-2.5, demonstrating that RL training with IXC-2.5-Reward improves instruction following and conversational ability without sacrificing performance in these areas.
@@ 5.1.2 Results on Reward Bench and RM-Bench
To prevent the chat model from generating overly long responses to artificially inflate rewards, we introduce length constraints on the ratio of chosen to rejected responses during training reward model IXC-2.5-Reward. The ablation study results of length constraints are present in Tab. 7. On the WildVision benchmark, we compute the average token length of the model’s responses. We observe a substantial increase in average token length, from 274 to 361, when length constraints were not applied. Surprisingly, removing length constraints yielded substantial improvements in open-ended benchmarks, achieving state-of-the-art results. Such observation is because these benchmarks do not penalize length in their evaluation prompts, judge models (e.g., GPT-4) tend to favor longer responses, even if they contain unnecessary details that detract from the user experience. As our focus is on optimizing user experience, not benchmark scores, we retain the length constraints. Following the precedent set by language-only benchmarks (e.g., [23]), we believe multi-modal Chat Arena and dialogue benchmarks should also address potential length and style biases in their evaluation protocols in future work.
@@
6 Conclusion and Future Work
We present IXC-2.5-Reward, a multi-modal reward model that is capable of multi-modal RL training, test-time scaling and data cleaning. IXC-2.5-Reward supports text, image, and video inputs and is trained on diverse, high-quality preference data to ensure robust preference discrimination. Using IXC-2.5-Reward, we further trained IXC-2.5-Chat via RLHF techniques to optimize the multi-modal user chat experience, focusing on providing detailed explanations and in-depth answers.
We believe that exploring multi-modal reward models with on-policy reinforcement learning algorithms holds significant promise for future research. For instance, while IXC-2.5-Reward can score multi-image and video inputs, existing reward benchmarks and applying RL algorithms for video alignment remain scarce. Moreover, extending the development of multi-modal reward models beyond chat to other domains, such as process reward models (PRMs) for multi-modal reasoning, is a promising direction.

### v2

@@ 5 Experiments
Table 4: Evaluation results on Reward Bench Lambert et al. (2024b). We report the performance of selective representative language-only RMs and previous multi-modal generative RMs.
@@
InternLM2-7B-Reward Cai et al. (2024) InternLM2-7B 99.2 69.5 87.2 94.5 87.6
InternLM2-20B-Reward Cai et al. (2024) InternLM2-20B 98.9 76.5 89.5 95.8 90.2
Skyword-Reward-Llama3.1-8B Liu et al. (2024a) Llama3.1-8B 95.8 87.3 90.8 96.2 92.5
INF-ORM-Llama3.1-70B Yang et al. (2024a) Llama3.1-70B 96.6 91.0 93.6 99.1 95.1
@@
QWen2-VL-7B Wang et al. (2024d) QWen2-7B 96.6 57.0 73.9 94.3 83.8
LLaVA-Critic-8B Xiong et al. (2024) LLaMA3-7B 96.9 52.8 81.7 83.5 80.0
IXC-2.5-Reward (Ours) InternLM2-7B 90.8 83.8 87.8 90.0 88.6
Table 5: Evaluation results on RM-Bench Liu et al. (2024b). We classify reward models into three types: sequence classifiers (Seq.), generative models, and implicit DPO models. Performance is reported across four domains (Chat, Math, Code, Safety) and three difficulty levels (Easy, Normal, Hard), along with average scores.
@@
Tulu-2-dpo-13b Ivison et al. (2023) Implicit 66.4 51.4 51.8 85.4 86.9 66.7 37.7 63.8
InternLM2-7B-Reward Cai et al. (2024) Seq. 61.7 71.4 49.7 85.5 85.4 70.7 45.1 67.1
InternLM2-20B-Reward Cai et al. (2024) Seq. 63.1 66.8 56.7 86.5 82.6 71.6 50.7 68.3
Nemotron-4-340B-Reward Wang et al. (2024f) Generative 71.2 59.8 59.4 87.5 81.0 71.4 56.1 69.5
URM-LLaMa-3.1-8B Lou et al. (2024) Seq. 71.2 61.8 54.1 93.1 84.0 73.2 53.0 70.0
Skyword-Reward-Llama3.1-8B Liu et al. (2024a) Seq. 69.5 60.6 54.5 95.7 89.0 74.7 46.6 70.1
@@
Instruction WildVision(0617) Lu et al. (2024c) Open 89.2 Hurst et al. (2024) 67.3 Xiong et al. (2024) 37.5 74.6
Following MIA${}_{(\text{val})}$ Qian et al. (2024) Open 88.6 Hurst et al. (2024) 80.7 Wang et al. (2024d) 80.4 84.0
& Chat MM-MT${}_{(\text{val})}$ Agrawal et al. (2024) Open 7.72 Hurst et al. (2024) 5.45 Wang et al. (2024d) 3.85 5.70
MM-Vet v2${}_{(\text{0613})}$ Yu et al. (2023) Open 71.8 Anthropic (2024) 58.1 Chen et al. (2024d) 45.8 54.8
Knowledge MMBench${}_{(\text{v1.1})}$ Liu et al. (2025) MCQ 85.7 SenseTime (2024) 82.7 Lu et al. (2024b) 79.4 79.0
MMMU${}_{(\text{val})}$ Yue et al. (2024) MCQ 70.7 Hurst et al. (2024) 56.2 Chen et al. (2024d) 42.9 44.1
MMStar Chen et al. (2024b) MCQ 72.7 SenseTime (2024) 63.2 Chen et al. (2024d) 59.9 59.6
Reasoning MathVista${}_{(\text{mini})}$ Lu et al. (2023) VQA 78.4 SenseTime (2024) 66.5 Lu et al. (2024a) 63.7 63.4
MathVerse${}_{(\text{vision-only})}$ Zhang et al. (2025b) VQA 54.8 Google (2024) 26.6 Liu et al. (2024c) 16.2 19.0
MathVision${}_{(\text{full})}$ Wang et al. (2024c) VQA 43.6 Google (2024) 22.0 Liu et al. (2024c) 17.8 18.8
Text-Rich TextVQA${}_{(\text{val})}$ Singh et al. (2019) VQA 82.0 Megvii (2024) 78.5 Li et al. (2024a) 78.2 81.3
ChartQA${}_{(\text{test})}$ Masry et al. (2022) VQA 81.2 Megvii (2024) 82.4 Yao et al. (2024) 82.2 80.5
OCRBench Liu et al. (2024d) VQA 89.4 SenseTime (2024) 82.2 Chen et al. (2024d) 69.0 70.0
5.1 Evaluation Results of IXC-2.5-Reward
Benchmarks. To evaluate IXC-2.5-Reward, we use diverse reward model benchmarks: (1) VL-RewardBench Li et al. (2024b), encompassing 1250 multi-modal problems addressing general understanding, hallucination, and reasoning challenges; (2) Reward-Bench Lambert et al. (2024b), with 2985 language-only problems including chat, chat hard, safety and reasoning; and (3) RM-Bench Liu et al. (2024b), comprising 1237 language-only problems across chat, math, code, and safety. RM-Bench defines three tracks (easy, normal, hard) that evaluate the sensitivity of reward models to subtle content variations and style biases. While Reward-Bench and RM-Bench are designed for reward models of language-only LLMs, we evaluate IXC-2.5-Reward on these benchmarks to demonstrate that our multi-modal reward model maintains strong language capabilities despite also processing image and video inputs.
5.1.1 Results on VL-RewardBench
Main Results. Tab. 3 presents the evaluation results of various multi-modal RMs on the VL-RewardBench Li et al. (2024b). Unlike previous multi-modal generative reward models, our IXC-2.5-Reward is a discriminative model that predicts a scalar reward. Our proposed IXC-2.5-Reward model, despite being an open-source 7B parameter model, outperforms all other open-source models. Notably, IXC-2.5-Reward achieves the highest overall accuracy (65.8%) among open-source models and the highest Macro Accuracy (70.0%) among all models, indicating its superior performance in handling diverse tasks within the VL-RewardBench.
Strong Performance on General Problems. The results in Table 3 reveal that IXC-2.5-Reward achieves a significantly higher accuracy (84.7%) on general problems compared to other generative RMs. We found the reason is attributed to these problems presenting a considerable challenge, often leading to tied judgments in previous LVLMs, whereas IXC-2.5-Reward demonstrates a greater ability to make correct classifications with different scalar scores.
5.1.2 Results on Reward Bench and RM-Bench
Main Results. We argue that multi-modal RMs should preserve strong language processing abilities despite the incorporation of image and video data during training. Consequently, we evaluate the performance of multi-modal reward models, including IXC-2.5-Reward, on Reward Bench (Tab. 4) and RM-Bench (Tab. 5). The results demonstrate that IXC-2.5-Reward achieves considerable performance and surpasses other multi-modal models on this benchmark.
Sensitivity to Content and Style. Consistent with findings in Liu et al. (2024b), IXC-2.5-Reward demonstrates sensitivity to subtle content variations and style biases, an issue often overlooked in multi-modal research. We believe further research is needed to enhance the robustness of multi-modal reward models.
5.2 Evaluation Results of IXC-2.5-Chat
Benchmarks. We select four representative benchmarks for evaluating the instruction following and in-the-wild chatting abilities of LVLMs. (1) The WildVision bench Lu et al. (2024c) uses prompts collected from user submissions, reflecting real-world multimodal interactions. (2) MIA-bench Qian et al. (2024) that is specially designed to evaluate instruction following. (3) MM-MT Agrawal et al. (2024) which is an instruction-following benchmark for multi-modal models, exhibits a strong correlation with LMSys-Vision ELO ratings Chiang et al. (2024). (4) MM-Vet Yu et al. (2023) that evaluate LVLMs on complex tasks such as language generation. These datasets contain open-ended questions and the referenced answers, and evaluation is performed using an LLM-as-a-Judge Zheng et al. (2023) approach, which involves using a judge model like GPT-4o Hurst et al. (2024) to predict scores.
We also report the performance on other categories, such as MMBench Liu et al. (2025), MMMU Yue et al. (2024) and MMStar Chen et al. (2024b) for general knowledge, MathVerse Zhang et al. (2025b) and MathVision Wang et al. (2024c) for math reasoning, TextVQA Singh et al. (2019), ChartQA Masry et al. (2022) and OCRbench Liu et al. (2024d) for text-rich document understanding. These benchmarks utilize multiple-choice questions (MCQ) or visual question answering (VQA), where responses are limited to short keywords and evaluated based on string matching.
Results on Instruction Following & Chat. Tab. 6 shows that IXC-2.5-Chat outperforms previous SOTA models across multiple benchmarks (WildVision, MIA, and MM-MT), demonstrating significant improvements in multi-modal understanding with instruction following ability and providing more comprehensive information for in-the-wild chat scenarios.
Results on Other Categories. On other categories (Knowledge, Reasoning, and Text-Rich), IXC-2.5-Chat performs comparably to the supervised fine-tuned (SFT) model IXC-2.5, demonstrating that RL training with IXC-2.5-Reward improves instruction following and conversational ability without sacrificing performance in these areas.
6 Conclusion and Future Work
We present IXC-2.5-Reward, a multi-modal reward model that is capable of multi-modal RL training, test-time scaling, and data cleaning. Using IXC-2.5-Reward, we further trained IXC-2.5-Chat via RLHF techniques to optimize the multi-modal user chat experience, focusing on providing detailed explanations and in-depth answers. We believe that exploring multi-modal reward models with on-policy reinforcement learning algorithms holds significant promise for future research, such as exploring reward benchmarks and RL algorithms for video alignment.
7 Limitations
The limitation of our work stems from the composition of our training data, which is primarily sourced from English language corpora. This reliance on English-centric data potentially limits the multilingual capabilities of our reward model. The English language datasets may reflect specific cultural viewpoints and societal biases prevalent in English-speaking communities. Future research should consider the incorporation of multilingual datasets to mitigate these limitations and enhance the generalizability and fairness of the multi-modal reward model.
Appendix
Appendix A More Experimental Results
@@
To prevent the chat model from generating overly long responses to artificially inflate rewards, we introduce length constraints on the ratio of chosen to rejected responses during training reward model IXC-2.5-Reward. The ablation study results of length constraints are present in Tab. 7. On the WildVision benchmark, we compute the average token length of the model’s responses. We observe a substantial increase in average token length, from 274 to 361, when length constraints were not applied. Surprisingly, removing length constraints yielded substantial improvements in open-ended benchmarks, achieving state-of-the-art results. Such observation is because these benchmarks do not penalize length in their evaluation prompts, judge models (e.g., GPT-4) tend to favor longer responses, even if they contain unnecessary details that detract from the user experience. As our focus is on optimizing user experience, not benchmark scores, we retain the length constraints. Following the precedent set by language-only benchmarks (e.g., Dubois et al. (2024)), we believe multi-modal Chat Arena and dialogue benchmarks should also address potential length and style biases in their evaluation protocols in future work.